# Leave No One Behind: Enhancing Diversity While Maintaining Accuracy in Social Recommendation

### v1

@@ 4.3 Diversity-Knowledge Transfer
Knowledge distillation has been actively studied for model compression in various fields  [27, 10, 39]. It transfers the knowledge captured by a teacher model through a large capacity into a student model, maintaining comparable performance. Different from general methods which transfer individual outputs from a teacher model to a student model point-wise, relational knowledge distillation (RKD) [24] was introduced to transfer relations of the outputs structure-wise.
To enhance recommendation diversity, we employ the embedding similarity between users and their friends as a diversity indicator. Notably, this similarity operates at a structure-to-structure level, as depicted in the right panel of Figure 2. Therefore, we adopt relational knowledge distillation to transfer knowledge from the teacher model to the student model:
@@
The angle-wise distillation loss $\mathcal{L}_{D}$ transfers the relationship of training example embeddings by penalizing angular differences. As angles encapsulate higher-order properties compared to distances, they offer greater efficacy in transferring relational information, affording the student model increased flexibility.
Given the complexity and noise inherent in real-world social relationships, we adopt a pragmatic approach by refraining from computing the similarity between individual user-friend pairs $(a,b)$ in $G_{s}$. Instead, we compute the average vector of all his social neighbors $p_{af}=mean(p_{\{b\}}),b\in\mathcal{N}_{s}(a)$, thereby smoothing out feature representations and mitigating noise in the social network.
@@ 4.4 Model Training
The learning of the student model consists of two tasks: recommendation and knowledge distillation task. The overall objective is defined as:
@@
where $\mathcal{L}_{R}$ is the recommendation task loss function, $\mathcal{L}_{D}$ is the knowledge distillation loss function, and $\beta$ is a hyperparameter that governs the trade-off between the two objectives. A larger $\beta$ prioritizes the acquisition of diversity knowledge, whereas a lower value emphasizes accuracy. Upon acquiring the combined representations for all users and items within the student model, we can subsequently predict user $a$’s preference for item $i$: $\tilde{r}_{a,i}(\theta_{s})=(q_{i}^{s})^{T}p_{a}^{s}$.
@@ 5.1.1 Datasets
In order to be consistent with previous research  [35, 41], we conduct experiments on three benchmark datasets. Yelp [^1] https://www.yelp.com/ is a popular online location-based social network that allows users to share their experiences. Ciao [^2] http://www.cse.msu.edu/ tangjili/trust.html is a popular social networking website where users can rate items, write reviews, and add friends. Flickr [^3] https://www.flickr.com/ is an online image-based social sharing platform. The statistics of these datasets are shown in Table 2.
@@ 5.1.2 Metrics
In this work, we adopt four widely used metrics to measure accuracy and diversity. For accuracy, we use Recall@K and NDCG@K [41]. Recall denotes computing the fraction of relevant items out of all relevant items. NDCG is the normalized discounted cumulative gain, which gives more weight to highly ranked resources and incorporates different relevance levels through different gain values. To measure diversity, we utilize Coverage@K and Entropy@K, commonly applied in diversified recommendation [25, 43]. Coverage measures the extent to which a recommendation result encompasses diverse items from the entire item pool. Entropy assesses the uniformity of item probabilities within the recommendation results. Higher Coverage@K and Entropy@K mean higher diversity. To save space, we only report top-100 recommendation results, noting that similar conclusions hold for other top-N recommendations.
@@ 5.1.3 Implementation Details
Our experiments are conducted on NVIDIA V100 GPUs with 32GB memory. For all the methods, we refer to the hyperparameter ranges provided in their original papers and use grid search to find the optimal set of hyperparameters. We employ the Adam optimizer [14] with a gradient descent-based approach, initializing the learning rate at 0.001. The batch size is set as 2000. The embedding size is fixed as 64 and the $L_{2}$ regularization parameter $\lambda$ is 0.001. The coefficient $\beta$ for knowledge distillation is searched among {$42.0,1.0,0.5,...,1e-$ 4}. The experiments were conducted five times, and results reported as averages, reflect the model’s performance. Additionally, early stopping is utilized to alleviate the over-fitting problem.
@@ 5.2 Main Results with Various Backbone Models
As DivSR is model-agnostic, we evaluate its performance with representative social recommenders. TrustMF [36] employs matrix factorization (MF) to embed users into low-dimensional spaces. SocialMF [11] is a regularization-based social recommendation model that constrains users’ latent vectors to be close to those of their social neighbors. DiffNet [35] utilizes graph convolutional networks to capture dynamic social diffusion in social graphs. MHCN [41] enhances social recommendation through self-supervised learning (SSL) on motif-induced hypergraphs. DESIGN [31] performs statistical data analyses to obtain a deeper understanding of the social influence theory.
@@
We train the base social models and their DivSR counterparts on three datasets. The overall recommendation results are shown in Table 3.
Here, we observe that social recommendation methods usually reduce recommendation diversity while improving recommendation accuracy compared to their non-social variants (except DESIGN and MHCN on Flickr). For DESIGN and MHCN on Flickr, we guess that the tendency that socially connected users to have similar preferences is not obvious. They just utilize social relations as supplemental information for self-supervised learning. For accuracy, GNNs-based methods outperform MF-based methods (i.e. DiffNet vs. TrustMF), which can be attributed to GNNs’ strong capability in modeling graph data. SSL-enhanced methods prove to be more effective than methods without SSL (i.e. MHCN vs. DiffNet), highlighting the effectiveness of self-supervised learning. Regarding diversity, GNN-based methods tend to generate more diverse recommendations than MF-based methods (i.e. DESIGN vs. SocialMF), which could be credited to their ability to equip users/items with more neighbors.
DivSR promotes diversity without excessively sacrificing accuracy, and in some cases, even improving it compared to these social recommendation backbones. For example, DivSR boosts diversity from 23% to 33% on average across the three datasets towards SocialMF, with accuracy fluctuating by less than 1%. Compared to GCN-based backbones, DivSR achieves a more substantial improvement in diversity over MF-based backbones. Meanwhile, DivSR generally exhibits a higher coverage while maintaining recall score compared to social recommendation backbones, indicating a better accuracy-diversity trade-off. It’s noteworthy that DivSR may occasionally demonstrate a slightly lower diversity than Social RS (w/o social), but the accuracy of the latter is markedly lower. These findings underscore the effectiveness of the knowledge distillation module in enhancing recommendation diversity while preserving high accuracy.
@@ 5.3 Comparison with Diversified Models
We conduct experiments for comparison with several diversified methods. MMR [4] leverages marginal relevance to balance relevance and diversity. DPP [5] is an elegant probabilistic model that is widely used for diversified recommendation. DivMF [13] regularizes a score matrix of an MF model to maximize coverage of top-k recommendation lists. DGRec [38] designs the submodular function to select a diversified subset of neighbors to enhance diversity.
@@
In Figure 3, two re-ranking methods (MMR, DPP), promote more diverse recommendations with higher diversity. However, their recommendation accuracy declines significantly, indicating their failure to handle the accuracy-diversity dilemma. In contrast, DivSR occupies the most favorable position in the upper-right quadrant, demonstrating that DivSR achieves the optimal accuracy-diversity trade-off. When compared to DGRec, DivSR exhibits a substantial increase in diversity while making only a minor sacrifice in accuracy.
@@ 5.4 Teacher Model Choice Analysis
We design two different sources for the supervised signal to train the student model. Uns.: It removes the teacher model and directly minimizes the user-friend similarity in the student model, $\mathcal{L}_{D}=\sum_{(a,b)\in G_{s}}{\left\|\psi_{S}(a,b)\right\|}_{2}$, akin to an unsupervised training approach. DivSR(M): It selects the model with the highest diversity from five non-social models as the teacher model. Specifically, we utilize MHCN (w/o social) as the sole teacher model for Yelp, DESIGN (w/o social) for Ciao, and DiffNet (w/o social) for Flickr.
@@
The results in Table 4 illustrate the comparison of different teacher choices. Firstly, the unsupervised method demonstrates that optimizing user features for diversity through similarity yields effectiveness. Despite potentially resulting in decreased accuracy, the unsupervised method often enhances diversity. Secondly, utilizing the supervised signal provided by the teacher model to guide diversity knowledge transfer proves beneficial for overall performance. DivSR and DivSR (M) consistently achieve superior trade-off results compared to the unsupervised method in most cases. Thirdly, DivSR typically outperforms DivSR (M) as it consistently ranks within the top-2 in comparisons. These results of variations underscore that meticulously optimizing user-friend similarity can lead to improved accuracy-diversity performance. Moreover, selecting different teacher models can yield significant performance gains.
@@ 5.5 Effect of Knowledge Distillation
To gain deeper insights into the effect of the knowledge distillation mechanism in DivSR, we present a visualization of user embeddings in Figure 5. Using the Louvain algorithm [3], we detect communities within the Flickr dataset. Subsequently, we randomly sample 2,000 users from the top five communities and apply t-SNE [20] to visualize their embeddings.
We can observe that TrustMF (w/o social) generates relatively messy user representation distributions, lacking clear community aggregation behavior. Conversely, TrustMF (with social) exhibits pronounced community segregation, with users within the same community clustering closely while remaining isolated from users in other communities. This segregation can impede user development and content diversity. However, our model exhibits a certain degree of community aggregation while reducing the extent of community segregation, thereby making the communities less isolated from each other. This result demonstrates that the distillation mechanism has successfully learned low user similarity in non-social recommendation systems, showcasing the effectiveness of DivSR.
@@ 5.6 Parameter Sensitivity
In this part, we investigate the impacts of different values of $\beta$ on trade-off performance between accuracy and diversity. As illustrated in Figure 5, as $\beta$ decreases, the recall gradually increases, reaching its peak at a value of 0.1, and then declines until convergence. Conversely, diversity gradually decreases until convergence. When $\beta$ decreases, the model assigns less weight to diversity learning, resulting in a decrease in diversity and an increase in accuracy. When $\beta$ falls within an appropriate range (i.e. 0.05-0.1), the model achieves a good balance between accuracy and diversity, exhibiting high accuracy and high diversity simultaneously. Finally, as $\beta$ approaches very small values, the constraint on diversity learning almost disappears, causing the model to degrade into a social recommendation model, with both accuracy and diversity converging to the performance of the backbone model. The coefficient $\beta$ plays a pivotal role in balancing the main task and the additional knowledge distillation task.
@@ 6 CONCLUSION
In this work, we study the trade-off between recommendation accuracy and diversity in social recommendation models. We empirically observe that several existing social recommendation methods tend to reduce diversity while improving accuracy compared to their non-social variants. In response to this challenge, we propose DivSR, a simple and model-agnostic solution that can be easily deployed on existing social recommendation models. DivSR operates by employing relational knowledge distillation techniques, facilitating the transfer of high-diversity structured knowledge from their non-social counterparts to the social recommendation models. Our findings demonstrate that, when prioritizing diversity, DivSR outperforms existing social recommender systems. Notably, DivSR achieves a superior accuracy-diversity trade-off compared to several diversified models, showcasing its efficacy in addressing the inherent tensions between accuracy and diversity in social recommendation systems.

### v2

@@ 4.3 Diversity-Knowledge Transfer
Knowledge distillation has garnered significant attention for model compression across various domains  [27, 10, 39]. It facilitates the transfer of knowledge from a teacher model, which typically has a large capacity, to a student model, thereby preserving comparable performance. Unlike conventional methods that transfer individual outputs from the teacher to the student on a point-wise basis, relational knowledge distillation (RKD) [24] is introduced to transfer relational information at the structural level of the outputs.
To enhance recommendation diversity, we leverage the embedding similarity between users and their friends as a diversity indicator. Notably, this similarity operates at a structure-to-structure level, as depicted in Figure 2. Consequently, we employ relational knowledge distillation to transfer knowledge from the teacher model to the student model:
@@
The angle-wise distillation loss $\mathcal{L}_{D}$ facilitates the transfer of relational information between training example embeddings by penalizing angular discrepancies. Since angles encapsulate higher-order properties compared to distances, they provide a more effective means of transferring relational information, thereby endowing the student model with greater flexibility.
Given the inherent complexity and noise in real-world social relationships, we adopt a pragmatic approach by refraining from computing the similarity between individual user-friend pairs $(a,b)$ in $G_{s}$. Instead, we compute the average embedding vector of all social neighbors of a user $a$, denoted $p_{af}=mean(p_{\{b\}}),b\in\mathcal{N}_{s}(a)$, which smooths out feature representations and mitigates noise within the social network.
@@ 4.4 Model Training
The learning process of the student model involves two tasks: the recommendation task and the knowledge distillation task. The overall objective is defined as follows:
@@
where $\mathcal{L}_{R}$ represents the recommendation task loss function, $\mathcal{L}_{D}$ denotes the knowledge distillation loss function, and $\beta$ is a hyperparameter that controls the trade-off between the two objectives. A larger value of $\beta$ prioritizes the acquisition of diversity knowledge, while a smaller value emphasizes accuracy. After obtaining the combined representations for all users and items within the student model, we can predict user $a$’s preference for item $i$: $\tilde{r}_{a,i}(\theta_{s})=(q_{i}^{s})^{T}p_{a}^{s}$.
@@ 5.1.1 Datasets
In order to be consistent with previous research  [35, 41], we conduct experiments on three widely used benchmark datasets. Yelp [^1] https://www.yelp.com/ is a popular online location-based social network that allows users to share their experiences. Ciao [^2] http://www.cse.msu.edu/ tangjili/trust.html is a well-known social networking website where users can rate items, write reviews, and add friends. Flickr [^3] https://www.flickr.com/ is an online image-based social sharing platform. The statistics of these datasets are summarized in Table 2.
@@ 5.1.2 Metrics
In this work, we adopt four commonly used metrics to measure accuracy and diversity. For accuracy, we utilize Recall@K and NDCG@K [41]. Recall computes the fraction of relevant items identified out of all relevant items. NDCG (Normalized Discounted Cumulative Gain) places greater emphasis on higher-ranked resources and incorporates varying relevance levels through different gain values. To measure diversity, we use Coverage@K and Entropy@K, which are frequently applied in diversified recommendation task [25, 42]. Coverage measures the extent to which a recommendation set covers diverse items from the entire item pool. Entropy assesses the uniformity of item probabilities within the recommendation set. Higher Coverage@K and Entropy@K mean greater diversity. To save space, we only report top-100 recommendation results, noting that similar conclusions hold for other top-N recommendations.
@@ 5.1.3 Implementation Details
Our experiments are conducted on NVIDIA V100 GPUs with 32GB memory. For all methods, we refer to the hyperparameter ranges provided in their original papers and perform grid search to identify the optimal set of hyperparameters. We use the Adam optimizer [14] with a gradient descent-based approach, initializing the learning rate at 0.001. The batch size is set as 2000, and the embedding size is fixed as 64. The $L_{2}$ regularization parameter $\lambda$ is 0.001. The coefficient $\beta$ for knowledge distillation is searched within the range {$42.0,1.0,0.5,...,1e-$ 4}. Each experiment is conducted five times, and the reported results represent the average performance across these runs. Additionally, early stopping is employed to mitigate overfitting.
@@ 5.2 Main Results with Various Backbone Models
Since DivSR is model-agnostic, we evaluate its performance with several representative social recommender systems. TrustMF [36] employs matrix factorization (MF) to embed users into low-dimensional spaces. SocialMF [11] is a regularization-based social recommendation model that constrains users’ latent vectors to be close to those of their social neighbors. DiffNet [35] utilizes graph convolutional networks to capture dynamic social diffusion in social graphs. MHCN [41] enhances social recommendation through self-supervised learning (SSL) on motif-induced hypergraphs. DESIGN [31] performs statistical data analyses to gain deeper insights into the theory of social influence.
@@
We train both the base social models and their DivSR counterparts on three datasets. The overall recommendation results are presented in Table 3.
From the results, we observe that social recommendation methods usually reduce recommendation diversity while improving recommendation accuracy compared to their non-social counterparts (except DESIGN and MHCN on the Flickr dataset). For DESIGN and MHCN on Flickr, we guess that the effect of social connections on user preferences is less pronounced. These models leverage social relationships as supplementary information for self-supervised learning, rather than directly modeling user preferences based on social influence. Regarding accuracy, GNNs-based methods consistently outperform MF-based methods (i.e. DiffNet vs. TrustMF), which can be attributed to the superior capability of GNNs in modeling graph data. SSL-enhanced methods prove to be more effective than methods without SSL (i.e. MHCN vs. DiffNet), highlighting the effectiveness of self-supervised learning in enhancing performance. In terms of diversity, GNN-based methods tend to generate more diverse recommendations compared to MF-based methods (i.e. DESIGN vs. SocialMF), likely due to their ability to incorporate a broader range of neighbors for users/items.
DivSR promotes diversity without significantly sacrificing accuracy and, in some cases, even improves it compared to these social recommendation backbones. For example, DivSR boosts diversity from 23% to 33% on average across the three datasets towards SocialMF, with accuracy fluctuating by less than 1%. When compared to GCN-based backbones, DivSR achieves a more notable improvement in diversity over MF-based backbones. Furthermore, DivSR typically demonstrates a higher coverage while maintaining a comparable recall score, indicating a better balance between accuracy and diversity. It’s noteworthy that DivSR may occasionally show slightly lower diversity than Social RS (w/o social), but the accuracy of the latter is significantly lower. These findings underscore the effectiveness of the knowledge distillation module in enhancing recommendation diversity while maintaining high accuracy.
@@ 5.3 Comparison with Diversified Models
We conduct experiments to compare DivSR with several diversified methods. MMR [4] leverages marginal relevance to balance relevance and diversity in recommendation lists. DPP [5] is a probabilistic model that is widely adopted for generating diversified recommendations. DivMF [13] regularizes the score matrix of an MF model to maximize the coverage of top-k recommendation lists. DGRec [38] designs a submodular function to select a diversified subset of neighbors, thereby enhancing diversity.
@@
As shown in Figure 3, two re-ranking methods, MMR and DPP, prioritize diversity by generating more varied recommendations. However, this comes at the cost of a significant drop in recommendation accuracy, highlighting their inability to effectively balance the accuracy-diversity trade-off. In contrast, DivSR positions itself in the upper-right quadrant, demonstrating its capability to achieve the optimal trade-off between accuracy and diversity. When compared to DGRec, DivSR offers a notable increase in diversity while incurring only a minor sacrifice in accuracy.
@@ 5.4 Teacher Model Choice Analysis
In this section, we explore two distinct approaches for providing the supervised signal to train the student model. Uns.: It removes the teacher model and directly minimizes the user-friend similarity in the student model, $\mathcal{L}_{D}=\sum_{(a,b)\in G_{s}}{\left\|\psi_{S}(a,b)\right\|}_{2}$, akin to an unsupervised training approach. DivSR(M): It selects the model with the highest diversity from five non-social models as the teacher model. Specifically, we use MHCN (w/o social) as the sole teacher model for Yelp, DESIGN (w/o social) for Ciao, and DiffNet (w/o social) for Flickr.
@@
The results presented in Table 4 compare the performance of different teacher model choices. Firstly, the unsupervised method demonstrates that optimizing user features for diversity by directly minimizing user-friend similarity proves effective. Although this may lead to a decrease in accuracy, it frequently results in improved diversity. Secondly, utilizing the supervised signal from the teacher model to guide the transfer of diversity knowledge yields notable benefits for overall performance. Both DivSR and DivSR (M) consistently achieve superior trade-off results compared to the unsupervised method in most cases. Finally, DivSR generally outperforms DivSR (M), consistently ranking among the top-2 in comparisons. These findings suggest that carefully optimizing user-friend similarity can significantly enhance the accuracy-diversity balance, and that selecting an appropriate teacher model can further enhance performance.
@@ 5.5 Effect of Knowledge Distillation
To gain a deeper understanding of the impact of the knowledge distillation mechanism in DivSR, we present a visualization of user embeddings in Figure 5. Using the Louvain algorithm [3], we detect communities within the Flickr dataset. We then randomly sample 2,000 users from the top five communities and apply t-SNE [20] to visualize their embeddings.
From the visualization, we observe that TrustMF (w/o social) produces relatively disorganized user representations, with no clear community aggregation. In contrast, TrustMF (with social) displays distinct community segregation, where users within the same community are closely grouped, while those in different communities remain well-separated. This pronounced segregation can limit user engagement and hinder content diversity. However, DivSR strikes a balance by exhibiting some degree of community aggregation while reducing excessive community segregation, thus fostering greater inter-community interaction. These results demonstrate that the distillation mechanism effectively encourages the student model to learn low similarity between users, as seen in non-social recommendation systems, underscoring the efficacy of DivSR.
@@ 5.6 Parameter Sensitivity
In this section, we investigate the impact of varying values of $\beta$ on the trade-off between accuracy and diversity. As shown in Figure 5, as $\beta$ decreases, recall gradually increases, reaching its peak at a value of 0.1, before declining toward convergence. In contrast, diversity steadily decreases until it stabilizes. As $\beta$ decreases, the model assigns less weight to diversity learning, leading to a reduction in diversity and an increase in accuracy. When $\beta$ is within an optimal range (i.e. 0.05-0.1), the model achieves a favorable balance, simultaneously maintaining high accuracy and high diversity. However, as $\beta$ approaches very small values, the constraint on diversity learning diminishes significantly, causing the model to degenerate into a purely social recommendation model, with both accuracy and diversity converging to the performance of the baseline model. The coefficient $\beta$ thus plays a critical role in balancing the primary recommendation task and the auxiliary knowledge distillation task.
@@ 6 CONCLUSION
In this work, we examine the trade-off between recommendation accuracy and diversity in social recommendation models. Through empirical analysis, we observe that many existing social recommendation methods tend to reduce diversity while improving accuracy compared to their non-social variants. To address this challenge, we propose DivSR, a simple yet model-agnostic framework that can be seamlessly integrated into existing social recommendation systems. DivSR leverages relational knowledge distillation techniques to transfer high-diversity structured knowledge from non-social models to social recommendation models. Our experiments demonstrate that DivSR significantly enhances diversity without substantially compromising accuracy. Moreover, DivSR achieves a superior accuracy-diversity trade-off compared to several diversified models, effectively mitigating the inherent tension between accuracy and diversity.